# LiDAR Remote Sensing Meets Weak Supervision: Concepts, Methods, and Perspectives

### v1

@@ 2.5 Summary and discussion
Weak supervision refers to training models using incomplete, impexact, inaccurate labels, or labels from other domains. This approach provides efficient and labor-saving support for LiDAR data interpretation. Current research mainly focuses on sparse and semi-supervised learning, with notable progress in semantic segmentation, achieving performance comparable to fully supervised models. However, progress in tasks such as object detection and instance segmentation has been more limited. Moreover, while there has been some research on imexact or inaccurate supervision methods, significant gaps remain, leading to the wastage of extensive labeled data (e.g., crowdsourced labels). With the development of foundation models, this issue is expected to be alleviated. Foundation models can automatically generate labels for LiDAR point cloud data, reducing reliance on manual annotation and significantly lowering labeling costs.
@@ 3 LiDAR Inversion with Weak Supervision
Figure 9: Workflow of LiDAR inversion with weak supervision. (a) canopy height estimation, (b) building height estimation, and (c) biomass estimation.
In LiDAR remote sensing, LiDAR point clouds are often combined with remote sensing images for large-scale inversion tasks (e.g., canopy height, building height, biomass). Optical imagery provides continuous surface coverage with high revisit frequency and rich spectral information, while LiDAR data serves as supervisory signals, capturing structural and elevation details. However, due to the sparsity of LiDAR data, it cannot cover all inversion areas, making the generation of large-scale continuous maps challenging. Field surveys, which collect data such as tree diameter, height, and leaf area index by setting sampling points and survey regions, can provide more supervision signals but are costly and resource-intensive. Additionally, regional differences and temporal variations can lead to domain shifts, which pose significant challenges to model accuracy and generalization. Weakly supervised learning is an effective approach to reduce costs and enhance model generalization. Current weakly supervised LiDAR inversion primarily includes generalized methods and using LiDAR as weak supervision signals. However, the application of other weak supervision techniques remains limited, and the methods are relatively simple.
In this section, we will review and summarize existing related work. We categorize the applications in LiDAR remote sensing into specific areas, including canopy height mapping, building height estimation, and biomass estimation, as shown in Fig. 9.
@@ 3.1 Canopy height estimation
Current methods for canopy height inversion primarily use remote sensing images as training data and point clouds as supervisory signals. One approach involves training in small regions and extrapolating to large-scale areas, as shown in Fig. 12. They train a model using drone lidar data as labels and GEDI/ICESat-2 lidar footprints with environmental auxiliaries as inputs, then apply the trained model across mainland China to produce a high-precision forest canopy height map. Another uses point clouds as weak supervisory signals to directly train on large-scale remote sensing images, yielding large-scale inversion results, as shown in Fig. 13. Their CNN model uses Sentinel-2 (S2) imagery and encoded geographic coordinates (latitude, longitude) as input, supervised by sparse GEDI LiDAR labels, to estimate dense canopy height and its predictive uncertainty (variance).
Extrapolation. In the early stages, airborne point clouds are often used as weak supervisory signals. For instance, Pourshamsi et al. [201] combine Polarimetric Synthetic Aperture Radar (PolSAR) parameters with limited airborne LiDAR data, using various machine learning algorithms to train on a subset of LiDAR samples and predict forest height over larger areas. The emergence of spaceborne LiDAR has brought new opportunities. [202, 203, 204, 197, 205, 206, 207, 208, 209] use ICESat-2 or GEDI data as supervisory signals and apply machine learning or deep learning to train on remote sensing images, ultimately mapping canopy height. For instance, Potapov et al. [204] integrate NASA GEDI LiDAR data with Landsat multispectral time series data and use a machine learning regression tree model to train a forest canopy height prediction model within the limited coverage of GEDI. They extend the model to a global scale, generating a 30-meter resolution global forest canopy height map for the year 2019. Liu et al. [197] integrate GEDI and ICESat-2 ATLAS data to map the spatially continuous distribution of forest canopy height in China. This method uses drone-lidar data as training samples, trains the model on small regional areas, and then applies the model across the entire country to generate a 30-meter resolution forest canopy height map. Wang et al. [207] combine GEDI LiDAR data with other remote sensing data to improve the estimation of wall-to-wall canopy height.
@@
Deep learning techniques have made groundbreaking progress in recent years in 2D image processing, providing new solutions for the analysis of remote sensing data. Many studies [14, 210, 211, 212] now focus on using deep learning to treat point clouds as weak supervision signals for training on large-scale data such as remote sensing images. For instance, Weber et al. [210] propose a unified deep learning-based model to predict global above-ground biomass density (AGBD), canopy height (CH), and canopy cover (CC) by fusing 13-channel inputs from Sentinel-1 and Sentinel-2 satellite imagery, multispectral data, and digital elevation models (DEMs). Fan et al. [211] utilize a customized PRFXception neural network incorporating receptive fields at different scales, which learns canopy height information from sparse LiDAR sampling points and makes continuous inferences over the entire image coverage area, resulting in canopy height estimation at 10 m resolution. Fayad et al. [212] propose a hybrid visual transformer model called Hy-TeC to generate high-resolution tree canopy height maps using multi-source remote sensing data (Sentinel-1 and Sentinel-2) as inputs and GEDI LiDAR data as sparsely supervised ground truth. The model employs an architecture that combines a transformer encoder with a convolutional decoder, combined with a dual loss function of classification and regression to enhance the accuracy of the height intervals.
Figure 14: The result of candidate ground or building photons extraction. The image is obtained from [213].
@@ 3.2 Building height estimation
Building height, as a key parameter of urban three-dimensional morphology, plays a critical role in energy consumption estimation, urban microclimate modeling, 3D reconstruction, population distribution analysis, and urban planning [214]. Traditional methods, such as field surveys, are accurate but time-consuming and labor-intensive, making them unsuitable for large-scale applications. Remote sensing provides a cost-effective alternative but faces challenges: airborne LiDAR data are expensive, SAR data struggle with single-building delineation, and shadow-based methods using high-resolution optical imagery, though practical, often rely on image metadata and building height annotations while ignoring azimuth variations, limiting their applicability. ICESat-2 satellite ATL03 photon data offer high-precision building height measurements but are restricted by orbital coverage, allowing measurement of only a limited number of buildings, as shown in Fig. 14.
@@
Current methods integrate ICESat-2 photon data with remote sensing imagery to leverage their combined strengths, overcoming limitations of traditional methods and improving accuracy and applicability in building height extraction. [215, 216, 217, 218, 219] train the model with small-scale data and generalize it to large-scale areas. For instance, Ma et al. [216] combine GEDI LiDAR data with features from Landsat-8, Sentinel-2, Sentinel-1, and terrain data to train random forest models across 68 subregions globally, generating a 150-meter resolution global urban building height map. Ma et al. [217] combine GEDI LiDAR, optical (Landsat-8, Sentinel-2), and radar (Sentinel-1) data to produce a 150-meter resolution building height map for the Yangtze River Delta using a random forest model to scale discrete GEDI samples to regional mapping. Kamath et al. [218] integrate LiDAR, ICESat-2, GEDI, and ALOS DSM data with machine learning to generate global building heights and urban canopy parameters (UCPs). Using LiDAR-derived nDSM and features like building area and population density, they train a Random Forest model to predict building heights, which is then applied to cities worldwide for urban climate simulations, thermal comfort assessments, and energy consumption predictions.
@@ 3.3 Biomass estimation
Current research [221, 222, 223, 224, 225, 226, 227, 228, 229, 230] often utilizes regression models, machine learning, or deep learning techniques combined with remote sensing imagery, airborne LiDAR, or satellite LiDAR data to train on small regions and then generalize to larger areas. Campbell et al. [221] combine ground measurements with airborne lidar (ALS) data to construct a local-scale random forest model for predicting aboveground biomass (AGB) within the ALS coverage. They then use the predictions of the local model as reference data and combine them with Landsat optical remote sensing data and GEDI lidar data to build a regional-scale model, enabling AGB estimation over a larger area. Jiang et al. [223] combine ICESat-2 lidar data and Sentinel-2 optical remote sensing data and use an optimized Extreme Learning Machine (ELM) algorithm to estimate the aboveground biomass (AGB) of natural forests in the eastern Tibetan Plateau. Vázquez-Alonso et al. [224] combine ground surveys, LiDAR data, and Sentinel-2 imagery with a random forest model to classify forest types and estimate aboveground biomass (AGB). Using LiDAR-derived features, they build regression models to generate a high-resolution AGB map for a 25 km² area. Duncanson et al. [225] use ground and airborne lidar data from 74 global forest sites to generate training data with simulated GEDI waveforms. They develop stratified regression models based on relative height (RH) metrics, optimized by geographic and plant functional type (PFT) stratification, variable selection, and data transformation. Deng et al. [226] use ICESat-2 ATL08 LiDAR and multi-source remote sensing data (Landsat 8, Sentinel-1/2, DEM) with a random forest model to map 30-meter resolution canopy height. They further estimate forest aboveground biomass and carbon storage in Shangri-La City using height-DBH models and biomass formulas. Oehmck et al. [229] use deep learning models like PointNet, KPConv, and Minkowski CNNs to estimate above-ground biomass (AGB) and wood volume directly from airborne LiDAR point clouds. Leveraging 3D structural data, this method outperforms traditional statistical models and is applied to nationwide data in Denmark to create high-resolution biomass maps.

### v2

@@ 3.6 Summary and discussion
Unlike structured and dense 2D imagery, LiDAR point clouds arise from active laser scanning, producing inherently sparse, irregular, and non-uniform measurements. This violates the dense spatial continuity underlying many weak supervision techniques such as label propagation or consistency regularization. Sparse far-field regions disrupt neighborhood-based propagation, leading to unstable supervision and error accumulation. Multi-echo data further entangle vertical semantics (e.g., canopy–understory–ground) and induce heterogeneous noise distributions, breaking the one-to-one label–instance assumption. Severe occlusions create irreversible “data voids” beyond what self-training can recover. In addition, non-stationary spatiotemporal variability introduces domain shifts and unstable noise fields, undermining stationary distribution assumptions widely adopted in WSL frameworks.
To address these modality-specific challenges, recent WSL methods integrate geometry- and noise-aware strategies: pseudo-labeling and multi-modal fusion compensate for supervision gaps; elevation-aware modules and geometric priors effectively mitigate semantic ambiguity in the vertical dimension; redesigned consistency regularization accommodates irregular sampling; shape priors handle occlusion-induced voids; and domain/test-time adaptation aligns features under platform shifts. However, most studies still focus on static cross-dataset settings and offer limited solutions to occlusion heterogeneity and dynamic spatiotemporal shifts—core characteristics of LiDAR remote sensing. Bridging this gap requires rethinking WSL as a semantic- and geometry-aware framework rather than a direct adaptation of image-based strategies.
In summary, WSL leverages weak signals through pseudo-labeling, multi-modal fusion, consistency regularization, and structural priors to alleviate LiDAR’s intrinsic challenges, with notable success in semantic segmentation. Yet progress on more complex tasks (e.g., detection, instance segmentation) remains limited, and key modality-specific issues such as waveform complexity, occlusion heterogeneity, and temporal drift are underexplored. The emergence of Foundation Models offers a promising path to address these gaps by generating high-quality labels at scale and reducing annotation costs.
Table 5: Open LiDAR-based inversion datasets. AGB denotes aboveground biomass, and CH denotes canopy height.
Dataset Purpose Time Extent Images Ground Truth
Scope Surface Sensor Res. Sensor Res.
[…]
Figure 9: Typical workflow of LiDAR-based inversion with weak supervision. (a) canopy height estimation, (b) building height estimation, (c) biomass estimation, and (d) water depth estimation.
@@ 4 LiDAR Inversion with Weak Supervision
In LiDAR remote sensing, LiDAR point clouds are often combined with remote sensing images for large-scale inversion tasks (e.g., canopy height, building height, aboveground biomass (AGB), water depth). Optical imagery provides continuous surface coverage with high revisit frequency and rich spectral information, while LiDAR data serves as supervisory signals, capturing structural and elevation details. However, due to the sparsity of LiDAR data, it cannot cover all inversion areas, making the generation of large-scale continuous maps challenging. Field surveys, which collect data such as tree diameter, height, by setting sampling points and survey regions, can provide more supervisory signals but are costly and resource-intensive. Additionally, regional differences and temporal variations can lead to domain shifts, which pose significant challenges to model accuracy and generalization. WSL is an effective approach to reduce costs and enhance model generalization. In LiDAR inversion tasks, the application of WSL primarily follows two technical approaches. The first is domain generalization (DG), where a model is trained in a source domain rich with dense LiDAR data (e.g., from ALS coverage) and is then generalized to unseen target domains. The second is incomplete supervision, which directly uses spatially sparse LiDAR data (e.g., from GEDI and ICESat-2) as supervisory signals to train a model end-to-end over a large target area. However, the application of incomplete supervision techniques remains limited Table 5 presents the commonly open datasets.
In this section, we will review and summarize existing related work. We categorize the applications in LiDAR remote sensing into specific areas, including canopy height mapping, building height estimation, biomass estimation, water depth estimation and other applications, as shown in Fig. 9.
@@ 4.1 Canopy height estimation
Current approaches for canopy height retrieval typically use remote sensing imagery as training data and point clouds for supervision. One method follows the domain generalization (DG) setting, where a model trained on a small area is extrapolated to unseen regions, as shown in Fig. 12. This approach uses ALS data as labels, with GEDI/ICESat-2 LiDAR footprints and environmental auxiliary variables as inputs, to train a generalizable model for producing large-scale, high-precision canopy height maps. Another method follows the incompletely supervised learning setting, which uses sparse point clouds as weak supervision to train a model directly on large-scale remote sensing imagery, as shown in Fig. 13. This approach often employs a Convolutional Neural Network (CNN) that takes Sentinel-2 (S2) imagery and encoded geographic coordinates as input. Supervised by sparse GEDI LiDAR labels, the model concurrently estimates dense canopy height and its prediction uncertainty (variance).
Domain generalization. In the early stages, ALS point clouds are often used as weak supervisory signals. For instance, Pourshamsi et al. [215] combine Polarimetric Synthetic Aperture Radar (PolSAR) parameters with limited ALS data, using various machine learning algorithms to train on a subset of LiDAR samples and predict forest height over larger areas. The emergence of spaceborne LiDAR has brought new opportunities. Some studies use ICESat-2 or GEDI data as supervisory signals and apply machine learning or deep learning to train on remote sensing images, ultimately mapping canopy height [216, 211, 217]. For instance, Potapov et al. [216] integrate NASA GEDI LiDAR data with Landsat multispectral time series data and use a machine learning regression tree model to train a forest canopy height prediction model within the limited coverage of GEDI. They extend the model to a global scale, generating a 30-meter resolution global forest canopy height map for the year 2019. Liu et al. [211] integrate GEDI and ICESat-2 ATLAS data to map the spatially continuous distribution of forest canopy height in China. This method uses ALS data as training samples, trains the model on small regional areas, and then applies the model across the entire country to generate a 30-meter resolution forest canopy height map. Wang et al. [218] combine GEDI LiDAR data with other remote sensing data to improve the estimation of wall-to-wall canopy height. Turubanova et al. [219] utilize bagged regression tree models, integrating high-precision LiDAR data with long-term Landsat time-series imagery, to address the challenge of accurately mapping annual tree canopy extent and height dynamics across the entire European continent from 2001 to 2021. Schwartz et al. [220] employ a U-Net model, using sparse but precise LiDAR height measurements from NASA’s GEDI mission as labels to train nationwide Sentinel-1 (radar) and Sentinel-2 (optical) imagery, and produce a 10-m-resolution forest canopy height map for France for 2020 that outperforms existing models in accuracy. Similarly, they [221] also train a U-Net model that generates high-resolution forest canopy height maps of the whole“Landes de Gascogne” forest. Dixon et al. [222] employ a spatio-temporal deep learning model (ST-CNN) to analyze time series of high-resolution PlanetScope imagery, supplemented by radar and solar irradiance data, and produce 3-m resolution annual forest canopy height maps. Zhang et al. [223] use an automated machine learning (AutoML) model to fuse multisource satellite imagery (Sentinel-1/2, Landsat 8, ALOS-2) with precise GEDI LiDAR measurements as labels to map high-accuracy, high-resolution tropical rainforest canopy height and aboveground biomass. Xu et al. [224] use precise but spatially discrete ICESat-2 LiDAR tree-height measurements as samples and train a Random Forest model to learn spectral and textural features from high-resolution GF-2 imagery, thereby producing citywide tree canopy height maps. ARFCNet [225] combines convolutional layers, self-attention, and upsampling mechanisms to efficiently integrate airborne imagery, radar, and digital elevation models, enabling ultra-high-resolution (1 m) mapping of urban forest canopy height. Su et al. [226] train two independent deep learning models and fuse their predictions to generate high-resolution annual tree canopy height maps covering 20 years (2004–2024) for Italy. Song et al. [227] integrate LiDAR data with multi-source satellite remote sensing features to train a deep neural network model (STHNN) and employ the Shapley Additive Explanations (SHAP) method for feature selection, ultimately achieving large-scale mapping across the entire Shenzhen municipality. Seppi et al. utilize [228] GEDI LiDAR data and SAOCOM radar imagery to train a machine learning model for generating high-precision forest canopy height maps. Jamaluddin et al. [229] propose a 3D spatio-temporal-spectral deep learning model that fuses Sentinel-1 time-series radar data with Sentinel-2 red-edge optical data to achieve high-precision estimation of canopy height over large-scale mangrove forests. Malambo et al. [230] integrate high-precision laser altimetry data from the ICESat-2 satellite with Landsat and other multi-source auxiliary data, employing gradient boosting regression trees (XGBoost), to generate the first high-precision 30 m resolution vegetation canopy height map for the contiguous United States. Wagner et al. [231] employ a U-Net model to analyze high-resolution aerial imagery and reference LiDAR data, generating a high-precision, sub-meter resolution tree canopy height map across the entire state of California. They also [232] train a U-Net model using high-resolution Planet satellite imagery and airborne LiDAR data to predict tree canopy heights across the entire Amazon rainforest. Fogel et al. [210] present Open-Canopy, a large-scale open benchmark dataset comprising SPOT satellite imagery from France and LiDAR height maps. Leveraging this dataset, they conduct comprehensive experiments and evaluations on several mainstream computer vision models—including U-Net and PVTv2—for the task of estimating forest canopy height from satellite imagery. Pauls et al. [10] achieve high-precision, high-resolution (10 m) global forest canopy height mapping by addressing key flaws in GEDI labels. They filter unreliable labels in steep terrain using SRTM data and train a U-Net on combined Sentinel imagery with a novel “shift-resilient loss function” to correct for systematic positional errors.
@@
Deep learning techniques have made groundbreaking progress in recent years in 2D imagery, providing new solutions for the analysis of remote sensing data. Many studies employ deep learning frameworks under an incompletely supervised learning paradigm, treating point clouds as a weak supervisory signal for training on large-scale remote sensing imagery [16, 233]. For instance, Weber et al. [233] propose a unified deep learning-based model to predict global AGB density, canopy height, and canopy cover by fusing 13-channel inputs from Sentinel-1 and Sentinel-2 satellite imagery, multispectral data, and digital elevation models. Fan et al. [234] utilize a customized PRFXception neural network incorporating receptive fields at different scales, which learns canopy height information from sparse LiDAR sampling points and makes continuous inferences over the entire image coverage area, resulting in canopy height estimation at 10 m resolution. Fayad et al. [235] propose a hybrid visual transformer model called Hy-TeC to generate high-resolution tree canopy height maps using multi-source remote sensing data (Sentinel-1 and Sentinel-2) as inputs and GEDI LiDAR data as sparsely supervised ground truth. The model employs an architecture that combines a transformer encoder, a convolutional decoder, and a dual loss function of classification and regression to enhance the accuracy of the height intervals. Pauls et al. [236] train a 3D U-Net model to process satellite image time series accounting for seasonal variations, using sparse spaceborne LiDAR (GEDI) data as weak supervision labels, thereby generating high-resolution, annually varying dynamic tree canopy height maps across the European continent. MARSNet [237] employs independent encoders tailored to sparse LiDAR, SAR, and optical imagery to extract modality-specific features, thereby enabling efficient fusion of these multi-source data to generate a continuous high-resolution forest canopy height map.
The principal approach to canopy height estimation leverages sparse, pointwise height measurements from spaceborne LiDAR missions (primarily GEDI and ICESat-2) as weak supervisory signals. Early studies and many baseline comparisons relied on traditional machine-learning algorithms such as Random Forests. Contemporary approaches predominantly employ convolutional neural networks, such as U-Net and ResNet, to learn spatial context from image patches, while recent work increasingly adopts Transformer-based architectures, often in hybrid CNN and Transformer configurations, to capture long-range dependencies and improve accuracy in complex scenarios, for example, tall forest canopies. Despite substantial technical progress, the literature faces several common challenges: (1) data-quality issues, including label noise and signal saturation in input imagery; (2) limited model transferability, with models trained in one region generalizing poorly to other geographic contexts and forest types; and (3) persistent resolution mismatches that complicate multi-source data fusion.
Figure 14: Map of Haidian District, Beijing, with the ground tracks of ICESat-2 and the locations of field plots used to verify the accuracy of building height estimation. The 1-meter resolution satellite image is sourced from Google Earth and serves as the base map. The image is obtained from [238].
@@ 4.2 Building height estimation
Building height, as a key parameter of urban 3D morphology, plays a critical role in energy consumption estimation, urban microclimate modeling, 3D reconstruction, population distribution analysis, and urban planning [239]. Traditional methods, such as field surveys, are accurate but time-consuming and labor-intensive, making them unsuitable for large-scale applications. Remote sensing provides a cost-effective alternative but faces challenges: ALS data are expensive, SAR data struggle with single-building delineation, and shadow-based methods using high-resolution optical imagery, though practical, often rely on image metadata and building height annotations while ignoring azimuth variations, limiting their applicability. ICESat-2 satellite ATL03 photon data offer high-precision building height measurements but are restricted by orbital coverage, allowing measurement of only a limited number of buildings, as shown in Fig. 14.
@@
Current methods integrate ICESat-2 photon data with remote sensing imagery to leverage their combined strengths, overcoming limitations of traditional methods and improving accuracy and applicability in building height extraction. Some studies adopt a domain generalization (DG) paradigm, training models on small-scale data and generalizing them to unseen large-scale areas [240, 241]. For instance, Zhou et al. [242] use precise in-situ building-height measurements from selected cities in North America and Europe as labels to train a physically-based regression model that interprets global Sentinel-1 SAR signals, thereby producing the world’s first high-resolution atlas of urban building heights. Ma et al. [241] combine GEDI LiDAR data with features from Landsat-8, Sentinel-2, Sentinel-1, and terrain data to train Random Forest models across 68 subregions globally, generating a 150-meter resolution global urban building height map. Ma et al. [9] combine GEDI LiDAR, optical (Landsat-8, Sentinel-2), and radar (Sentinel-1) data to produce a 150-meter resolution building height map for the Yangtze River Delta using a Random Forest model to scale discrete GEDI samples to regional mapping. Kamath et al. [243] integrate LiDAR, ICESat-2, GEDI, and ALOS DSM data with machine learning to generate global building heights and urban canopy parameters. Using LiDAR-derived nDSM and features like building area and population density, they train a Random Forest model to predict building heights, which is then applied to cities worldwide for urban climate simulations, thermal comfort assessments, and energy consumption predictions. Similarly, Tang et al. [244] combine ICESat-2 elevation measurements with Sentinel satellite observations and a digital surface model (DSM), and employ a Random Forest regression to map built-up heights in the two megacities, New York and Shenzhen. Ma et al. [245] address the challenges of global urban building height data lacking timeliness and high resolution by employing a Random Forest model. They utilize precise spaceborne LiDAR (GEDI) data as training labels and integrate global multi-source satellite imagery (Landsat, Sentinel, etc.) as features to generate a new 150 m resolution global urban building height map for 2020.
Currently, large-scale building height estimation research has converged on a prevailing technical paradigm that leverages spaceborne LiDAR data as the principal weak supervisory signal for label generation, integrated with machine learning models and multi-source remote sensing imagery to enable spatially continuous mapping. Specifically, training labels are primarily derived from two distinct spaceborne LiDAR systems. The first is GEDI, whose relatively large footprint (approximately 25 m) produces mixed relative height metrics that integrate information from the ground surface, vegetation, and built structures. The second is ICESat-2, whose denser photon data, when processed through advanced algorithms such as RANSAC, yield discrete yet more accurate estimates of building height. To support spatial extrapolation, models commonly incorporate time-series features derived from Sentinel-1 backscatter coefficients and Sentinel-2 optical imagery (for instance, shadow dynamics) as input variables. In terms of model selection, Random Forest remains predominant owing to its robustness in handling high-dimensional and heterogeneous data, while some studies employ Support Vector Regression (SVR) as an alternative.
Nevertheless, this paradigm faces intrinsic challenges related to the sparsity and quality of training samples. The linear sampling pattern of spaceborne LiDAR produces extremely sparse and spatially uneven observations, further affected by noise and geolocation errors that degrade label accuracy. This sparsity and imbalance, particularly the underrepresentation of tall buildings, often lead to a “saturation effect,” causing systematic underestimation of building heights exceeding approximately 30 m. Furthermore, models trained on sparse orbital tracks within a limited region struggle to capture diverse urban morphologies, thereby restricting spatial transferability and generalization. Finally, resolution mismatches in fusing multi-source and multi-scale remote sensing data (for example, 10 m imagery combined with 150 m grid data) constitute a major technical bottleneck for achieving high-precision building height estimation.
@@ 4.3 Biomass estimation
Current research typically employs regression models, machine learning, or deep learning techniques in conjunction with remote sensing imagery, airborne LiDAR (ALS), or satellite LiDAR data. Within a domain generalization (DG) paradigm, models are trained on small-scale regions and subsequently generalized to unseen spatial extents [246, 247, 248]. Campbell et al. [246] combine ground measurements with ALS data to construct a local-scale Random Forest model for predicting AGB within the ALS coverage. They then use the predictions of the local model as reference data and combine them with Landsat optical remote sensing data and GEDI LiDAR data to build a regional-scale model, enabling AGB estimation over a larger area. Jiang et al. [249] combine ICESat-2 LiDAR data and Sentinel-2 optical remote sensing data and use an optimized Extreme Learning Machine (ELM) algorithm to estimate the AGB of natural forests in the eastern Tibetan Plateau. Vázquez-Alonso et al. [12] combine ground surveys, LiDAR data, and Sentinel-2 imagery with a Random Forest model to classify forest types and estimate AGB. Using LiDAR-derived features, they build regression models to generate a high-resolution AGB map for a 25 km² area. Duncanson et al. [250] use ground and ALS data from 74 global forest sites to generate training data with simulated GEDI waveforms. They develop stratified regression models based on relative height metrics, optimized by geographic and plant functional type (PFT) stratification, variable selection, and data transformation. Deng et al. [251] use ICESat-2 ATL08 LiDAR and multi-source remote sensing data (Landsat 8, Sentinel-1/2, DEM) with a Random Forest model to map 30-meter resolution canopy height. They further estimate forest AGB and carbon storage in Shangri-La City using height-DBH models and biomass formulas. Oehmck et al. [248] use deep learning models like PointNet [33], KPConv [36], and Minkowski [39] to estimate AGB and wood volume directly from ALS point clouds. Leveraging 3D structural data, this method outperforms traditional statistical models and is applied to nationwide data in Denmark to create high-resolution biomass maps. Nascetti et al. [208] train deep-learning models (e.g., U-Net and Transformer) using multimodal satellite time-series data (Sentinel-1 radar and Sentinel-2 optical imagery) as inputs and high-accuracy airborne LiDAR-derived aboveground biomass (AGB) as labels, thereby addressing the challenge of estimating forest biomass with high accuracy and fine spatial resolution. Liu et al. [252] combine high-resolution PlanetScope satellite imagery with airborne LiDAR data to train a U-Net deep-learning model and integrate allometric growth equations, thereby addressing the prior inability to accurately quantify carbon stocks of trees outside forests across Europe. Guo et al. [253] fuse Sentinel satellite imagery to extend discrete GEDI LiDAR height points into a continuous forest canopy height map, which is further used to construct high-precision forest aboveground biomass (AGB). Ma et al. [254]first integrate forest inventory data with ICESat-2 LiDAR data to generate high-precision aboveground biomass (AGB) sample points, then use these points as training labels to learn their relationships with wall-to-wall Landsat-8 imagery, topographic, and meteorological data via machine learning models, ultimately producing a continuous AGB distribution map across the entire study area. Schwartz et al. [255] utilize a deep learning model, with GEDI data as labels, to train the model to directly learn from single-pass Sentinel-1 radar and Sentinel-2 optical satellite imagery and estimate forest height maps. Bueno et al. [256] utilize GEDI LiDAR biomass data as labels to train a machine learning model for generating a complete forest biomass map across the entire hurricane-impacted area. Su et al. [257] first utilize U-Net to generate high-resolution canopy height maps, which are then input into a Random Forest model integrated with ground-measured data to ultimately estimate and map aboveground biomass. Zurqani et al. [258] utilize the Google Earth Engine platform and select the gradient boosting tree (GBT) algorithm to fuse GEDI LiDAR, multi-source satellite imagery, and terrain data for estimating forest aboveground biomass. Zhang et al. [259] fuse ICESat-2 LiDAR data to obtain optimal forest canopy height, which is incorporated as a key feature into a machine learning model, thereby significantly enhancing the accuracy of forest aboveground biomass estimation. Cui et al. [260] fuse GEDI with multi-source satellite remote sensing data (Sentinel-1 and Sentinel-2) using machine learning to provide more precise and detailed estimates of biomass consumption and carbon emissions for the 2020 Beachie Creek Fire than existing global products. Contreras et al. [261] successfully achieve precise estimation and mapping of aboveground biomass in large-scale olive orchards across the Mediterranean by integrating multi-source remote sensing data—including GEDI LiDAR, optical, and radar data—and employing machine learning models. Cai et al. [262] use deep learning models integrated with multi-source remote sensing data to produce, for the first time, annual 30 m resolution maps of forest aboveground biomass across China from 1985 to 2023. Considering that traditional biomass estimation models often overlook spatial autocorrelation (where nearby features are more similar) and heterogeneity (where relationships or patterns vary with spatial location) in geospatial contexts, Min et al. [263] propose a novel machine learning model that integrates spatial correlation and heterogeneity. By combining sparse GEDI LiDAR data with continuous multi-source remote sensing imagery, they successfully map high-precision forest aboveground biomass across the western United States.